OWL shift summary for Dec 28, 1999

ugur@cuhep.Colorado.EDU
Tue, 28 Dec 1999 06:13:36 -0700 (MST)

Shift Summary for OWL shift, Dec 28, 1999
========================================================================
On Shift: R.Kessler, U,Erkarslan
========================================================================
Hours of beam available at all intensities: 6.5 hr
Hours of beam available at nominal intensity: 6.5 hr
Hours of E832 data (special runs not included): 5.8 hr
Hours of special runs (laser scans, muon runs...): --
========================================================================
Accounting of lost time:lost CSI quadrant, drt_fresh_start
========================================================================
Runs ended this shift:
Run 15310 - BEAM_799 - in progress

========================================================================
Tapes brought to FCC this shift: UPA268-272
UPB268-272
UPC268-272
UPS291-292
========================================================================
Accesses during shift: None
========================================================================
Overview of shift:
E799 physics data at 9E12 with Many masseges from pipline.
[status show TEMP alarm on rack 1]

========================================================================
Description of Shift:

The only excitement was when the scaler alarms went off for
L1-B10/SEM and L1-B11/SEM ratios (hyperon triggers). Each of
these ratios dropped by a factor of 2 for several spills in run 15310;
the other 799 kaon trigger/SEM ratios did NOT change. See L1/SEM ratio
vs. spill plots on p.145, Logbook XXIII.
I suspect DCX-hole-ORS since they are unique to the hyperon triggers.
Anyway, to stop the anoying alarms I reduced the low limit by
about 5%.

We got a few spills above 1E13, so we called MCR to reduce it
down to 9.5E12.


01:55 split tape request(UPF067).

lost CSI quadrant. ECC tries to reset pipline. But it didn't slove the
problem. He decided to stop the run and drt_fresh_start. Change the
tapes. now, everthing is ok


========================================================================
Follow up Notes on Previous Shift:

========================================================================
Items to Follow up on Next Shift:


========================================================================